OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

Elimination: 2x + 5y = -3 + -15x - 5y = -10 = -13x = -13 /-13 x = 1

OpenStudy (anonymous):

Now to find y, you just have to substitute the 1 for x in the equation: 2x + 5y = -3 2(1) + 5y = -3 2 + 5y = -3 -2 5y = -5 /5 y = -1

OpenStudy (anonymous):

The elimination method is easier in this situation because the -5y and +5y from both equations already cancel out when the equations are added together. Your final solution would be: x = 1 y = -1 so, (1,-1) Hope I helped!

OpenStudy (anonymous):

Oh my gosh yes , thank you.!

OpenStudy (anonymous):

No problem:)